OpenStudy (anonymous):

A particle, initially at rest, moves along the x-axis such that its acceleration at time t>0 is given by a(t)=cos(t). At the time t=0, its position is x=3. A) Find the velocity and position functions for the particle. B) Find the values of t for which the particle is at rest.

OpenStudy (anonymous):

We will start with velocity.

OpenStudy (anonymous):

Alright :D

OpenStudy (anonymous):

\[ v(t)-v(0) = \int_0^t\cos(t)~dt \]

OpenStudy (anonymous):

They have already given us that \(v(0) = 0\), so that simplifies things for us.

OpenStudy (anonymous):

Oh so "initially at rest" means velocity right

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Can you integrate?

OpenStudy (anonymous):

Uhm That would mean that the acceleration would be sin(t) + C

OpenStudy (anonymous):

No, there is no constant of integration here, because it isn't an indefinite integral.

OpenStudy (anonymous):

We used a definite integral, so any constant of integration was canceled out through subtraction.

OpenStudy (anonymous):

I don't think I've learned definite/indefinite yet

OpenStudy (anonymous):

Definite integrals are ones where we have limits of integration. In this case our limits were \(0\) to \(t\). Indefinite ones don't have limits of integration.

OpenStudy (anonymous):

Actually I think I've only learned indefinite

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Technically, I should have written:\[ v(t)-v(0) = \int_0^t\cos(x)~dx \]Or something, because your variable of integration can't be a limit of integration as well, but is isn't a bit issue.

OpenStudy (anonymous):

Anyway, we should get: \[ v(t)-v(0) = \sin(t)\\ v(t)-0 = \sin(t)\\ v(t) = \sin(t) \]

OpenStudy (anonymous):

That is our velocity. The next thing to find is our position.\[ x(t)-x(0) = \int_0^t \sin(u)~du \]

OpenStudy (anonymous):

So I'm on the first chapter of integration, so I'm not very familiar with the vocab. but why don't we include C into the answer.

OpenStudy (anonymous):

Did you learn the fundamental theorem of calculus?

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

Okay, suppose you have a function \(f(x)\) with a derivative \(f'(x)\).

OpenStudy (anonymous):

One of the things the fundamental theorem of calculus says is that: \[ \int_a^bf'(t)~dt = f(b)-f(a) \]

OpenStudy (anonymous):

I also haven't learned the little notation on the ends of the integral sign either :c

OpenStudy (anonymous):

So you really haven't learned definite integrals?

OpenStudy (anonymous):

Nope, only indefinite

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

Okay, let's just suppose then that: \[ v(t) = \sin(t) + C \]

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

In that case we need to solve for \(C\) by using \(v(0)=0\).

OpenStudy (anonymous):

C=0

OpenStudy (anonymous):

So that would make v(t) = sin(t)

OpenStudy (anonymous):

and the position function is x(t)= -cos(t)

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

The integral of the Velocity function would be Position right

OpenStudy (anonymous):

But how did you find \(C\)?

OpenStudy (anonymous):

well you said that v(0) = 0 so i plugged it into the V(t) = sint +c 0=sin(0) +C C=0